# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

### v1

@@ Abstract
We present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential of LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
Currently, NOAA’s publicly available records on their website are available for 2000-2025. We expected earlier records to be available and submitted a data request to NOAA for documents covering 1972–1999, but have not received a response to date. As a result, the dataset currently covers 2000-2025. If older cloud seeding reports from 1972–1999 are recovered and made publicly available, the same pipeline can be applied with minimal modifications to expand the dataset’s coverage.
@@
To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2020 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest.
@@
For example, while there have been numerous recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2017 SNOWIE campaign in Western Idaho[3]), there has been little qualitative analysis of reported historical operational cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change, this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification usage[4].
@@ Methods
Figure 1: NOAA Data Extraction Pipeline using LLM
@@
After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [9]. We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing. All prompt versions and evaluation scripts are available in the GitHub repository [8].
@@
Between 2000 and 2025, most U.S. cloud seeding activity occurred in Utah, Colorado, and eastern California, primarily aimed at enhancing snowpack through ground-based silver iodide dispersion. This pattern is consistent with the nine western states known to have operational cloud seeding programs: California, Colorado, Idaho, Nevada, New Mexico, North Dakota, Texas, Utah, and Wyoming [4]. In Texas, by contrast, efforts focus on increasing rainfall and suppressing hail using a combination of airborne silver iodide dispersion and ground-based shockwave generation. Overall, cloud seeding activity declined between 2002 and 2023, but showed an uptick in 2024, with 2025 also on track for increased activity.
Figure 3: U.S. Cloud Seeding Activity by State (2000-2025)
Figure 4: Cloud Seeding Activity over Time (2000-2025)
Figure 5: Purpose of Cloud Seeding Activity
@@
A key challenge in using LLMs for data extraction is their generative nature. Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material. In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
@@
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-ground-truth.py.
Table 1: Field-Level Accuracy (n=200)
@@
We experimented with various OpenAI models to improve accuracy, including: gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3. This allowed us to observe how accuracy changed with model complexity and specialization. Ultimately, OpenAI’s ”more affordable reasoning model”, o4-mini, achieved the best balance of performance and cost. When comparing the different model variants, the same prompt was used (Prompt C).
Table 2: Model Comparison (n=50) using Prompt C
LLM Accuracy (%)
gpt-4.1 88.33%
gpt-4.1-mini 83.67%
gpt-4o-mini 86.67%
o3 90.33%
o4-mini 88.00%
@@
We also evaluated how different prompts influenced extraction accuracy, holding the model constant with gpt-4.1-mini. We compared three prompt formats: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and chain-of-thought (Prompt C) that encouraged step-by-step reasoning. Prompt C, using chain-of-thought reasoning, substantially outperformed the others, achieving an accuracy of 87.67%. This suggests that explicit reasoning instructions with examples can improve structured data extraction accuracy.
Table 3: Prompt Comparison (n=50) using gpt-4.1-mini
Prompt Accuracy (%)
Prompt A (Detailed) 72.00%
Prompt B (Concise) 57.00%
Prompt C (Chain-of-Thought) 87.67%
Together, these validation efforts demonstrate that LLM-based data extraction can be improved with careful model selection and prompt design. However, while we tested several variants, we did not exhaustively test every permutation of models-and-prompts listed. Further experimentation might reveal additional insights.
@@
All weather modification reports are self-reported by project operators and may not represent a complete record of weather modification activity (e.g., if there are omissions). For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day. Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [10].
@@
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor. Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.

### v2

@@ Abstract
Cloud seeding, a weather modification technique aimed at locally enhancing precipitation, has been used in the Western United States since the 1940’s. However, datasets to facilitate large-scale analysis of cloud seeding efforts are not currently available. To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential for LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2000 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[3].
@@
For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[4]), there has been little qualitative analysis of reported historical operational cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[5], this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of large-scale weather modification[3].
@@ Methods
Figure 1: NOAA Data Extraction Pipeline
@@
After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [10]. We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing (shown below). Chain-of-thought reasoning guides the model to break down complex tasks into intermediate steps, which improves performance on tasks requiring multi-step inference [11].
All prompt versions and evaluation scripts are available in the GitHub repository [9].
@@
Present your final extracted results concisely as follows, in lowercase, comma-separating multiple values when applicable.
Do not include commentary, explanations, or placeholders. Leave fields blank if truly unknowable after exhausting all inference methods.
@@
Between 2000 and 2025, cloud seeding activity in the United States was geographically concentrated in a few key states (Figure 3). California, Colorado, Utah, and Texas accounted for the majority of operations, with activity highest in western states that rely on snowpack for water supply. These records align with states that maintain active cloud seeding programs.
Over time, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4). This timeline shows that while cloud seeding has remained a consistent tool, its usage has fluctuated, likely due to policy changes, drought conditions, and funding cycles.
The primary stated purpose of cloud seeding events was to increase snowpack, followed by increasing precipitation and rain (Figure 5). Snowpack enhancement is particularly important in mountainous regions where runoff from melting snow supplies reservoirs and irrigation systems.
Finally, the types of agents and apparatus used vary significantly (Figure 6). Silver iodide is by far the most common seeding agent, especially in ground-based and airborne operations. Other agents such as calcium chloride and carbon dioxide are used less frequently. Ground deployment is the most prevalent method, particularly when using silver iodide, although a mix of airborne and dual (ground+airborne) approaches is also present.
Figure 3: U.S. Cloud Seeding Activity by State (2000–2025). States with active weather modification programs show the highest number of recorded operations.
Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
Figure 5: Purpose of Cloud Seeding Activity (2000–2025). Snowpack enhancement is the leading purpose, followed by increasing precipitation and rain.
Figure 6: Agent and Apparatus Used for Cloud Seeding (2000–2025). Silver iodide dominates among agents and is deployed most frequently using ground-based apparatus.
@@
A key challenge in using LLMs for data extraction is their generative nature. Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material [12]. In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
@@
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py.
Table 1: Field-Level Accuracy (n=200) using o4-mini LLM
@@
We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10]. This comparison allowed us to assess how accuracy varied with model complexity and specialization. Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document. While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
Each accuracy value in Table 2 represents the average field-level accuracy across 50 annotated NOAA reports, using a consistent prompt (Prompt C). For each report, model outputs were compared to human-labeled values across all fields, and the proportion of correct fields was averaged across the dataset. Although large language models are known to generate outputs probabilistically, we observed stable performance across repeated runs of the same model.
Table 2: Model Comparison (n=50)
LLM Overall Average Accuracy (%)
gpt-4.1 93.33%
gpt-4.1-mini 93.33%
gpt-4o-mini 91.67%
o3 96.33%
o4-mini 95.00%
@@
We also evaluated how different prompts influenced extraction accuracy, holding the model constant (o4-mini). Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning. Prompt C achieved the highest accuracy at 95.00%, suggesting that explicit reasoning steps can substantially improve structured data extraction.
Prompt B performed poorly (44.00%) due to frequent missing values in the outputs, likely caused by its lack of field-level guidance and examples. In contrast, Prompt A performed better (93.67%) and benefitted from including worked examples for each field. While not explicitly labeled as chain-of-thought, Prompt A was influenced by the same principles, demonstrating that even partial integration of reasoning structure can enhance accuracy.
Table 3: Prompt Comparison (n=50) using o4-mini LLM
Prompt Overall Average Accuracy (%)
Prompt A (Detailed) 93.67%
Prompt B (Concise) 44.00%
Prompt C (Chain-of-Thought) 95.00%
Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering. Chain-of-thought reasoning and example-driven prompts, in particular, are well-suited for scientific data extraction from complex archival documents. While our evaluation covered key variants, further experimentation could uncover additional insights.
@@
All weather modification reports in this dataset are self-reported by project operators and may not represent a complete record of U.S. weather modification activity, particularly in cases of unsubmitted or omitted filings.
As of this writing, NOAA’s publicly available records span the years 2000–2025. We anticipated earlier records to be accessible and have submitted a formal data request to NOAA for archived documents covering 1972–1999. If these older reports are recovered and released, the current extraction pipeline can be applied with minimal modification to extend the dataset’s temporal coverage.
For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day. Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [13].
@@
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor. Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/US-Cloud-Seeding-Analysis.